OpenStudy (anonymous):

Determine the length(distance) of the line segment with endpoints (4,-3) and (-1,10)

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

hartnn (hartnn):

could u find the distance ?

OpenStudy (anonymous):

√(4-(-1))^2 +(-3-10)^2 =√(4+1)^2 + (-13)^2

hartnn (hartnn):

thats correct, go ahead.

OpenStudy (anonymous):

√(5)^2 + (-13)^2 5+(-13) =-8? or √(25) + (169) =√194

hartnn (hartnn):

absolutely correct!

OpenStudy (anonymous):

Which one will it be? the √194 or -8, pls

hartnn (hartnn):

\(\huge \sqrt{194}\)

hartnn (hartnn):

got it ?

OpenStudy (anonymous):

yep

hartnn (hartnn):

good :)